# Dinkel: Testing Graph Database Engines via State-Aware Query Generation

### v2

@@ 5.5. Comparison
Bug Latency Study. To demonstrate that Dinkel can find bugs missed by existing approaches (Kamm et al., 2023; Hua et al., 2023; Jiang et al., 2023b; Zhuang et al., 2023; Mang et al., 2024), we investigate the latencies of the bugs found by Dinkel. For each bug, we check whether its bug-inducing commit was created before the years when existing approaches were published. If Dinkel finds some long-latent bugs, we can conclude that Dinkel can find bugs missed by existing approaches. This comparison is reasonable and objective because: (1) none of the bugs found by Dinkel are marked as duplicated by GDBMS developers, which means that no approach found these bugs until Dinkel found them; and (2) all existing approaches have extensively tested Neo4j and RedisGraph (Kamm et al., 2023; Hua et al., 2023; Jiang et al., 2023b; Zhuang et al., 2023; Mang et al., 2024), which means that in these two GDBMSs, no approach found the long-latent bugs found by Dinkel during their evaluation.
GDsmith (Hua et al., 2023), GDBMeter (Kamm et al., 2023), GraphGenie (Jiang et al., 2023b), and GAMERA (Zhuang et al., 2023) were published in 2023, and GRev (Mang et al., 2024) was published in 2024. Therefore, for each bug, we checked whether its bug-inducing commit was created before 2024 (i.e., in 2023 or earlier) and 2023. Table 10 shows the results. Among the 49 bugs Dinkel found in Neo4j and RedisGraph, 33 bugs already existed before 2023, indicating that all the existing approaches proposed after 2023 failed to find these 33 bugs in their extensive evaluation. In addition, 14 bugs were introduced to Neo4j and RedisGraph between 2023 and 2024, while GRev, which was proposed after 2024, cannot find these bugs. These results indicate that existing approaches indeed miss many long-latent bugs, while Dinkel can effectively find them.
Table 10. Latency of the logic bugs found by Dinkel
@@
Neo4j 32 30 20 2016
RedisGraph 17 17 13 2020
Apache AGE 11 11 7 2021
Total 60 58 40 2016
We analyze the bug-triggering queries of these bugs and conclude the reason why existing approaches missed these bugs as the following: (1) the bug-triggering queries (e.g., the queries shown in Figure 3 and Figure 6) contain data dependencies, which existing approaches lack support for handling; and (2) the bug-triggering queries contain Cypher features (e.g., CALL and FOREACH in the query shown in Figure 16) that are not supported by existing approaches. The following is an example of the 33 bugs.
1WITH [] AS n0 ORDER BY null
2CALL {
3 WITH [] AS n1 ORDER BY null
4 UNWIND [0] AS x
5 UNWIND [x] AS n2
6 RETURN n2 AS n3
7}
8FOREACH (n4 IN null | MERGE ({key:n3}))
Figure 16. A query triggering a Neo4j internal error—Neo4j-
Error: ExecutionFailed (arraycopy: last destination index 6 out of bounds for object array[5]).
Example: Query using data dependencies and advanced Cypher features. Figure 16 shows a bug in Neo4j, which is present in 2022 but none of the existing approaches found it. The bug-triggering query involves complicated data dependencies and advanced Cypher clauses (e.g., FOREACH and UNWIND). This query constructs an empty array [] and invokes a subquery using the CALL clause. The subquery also constructs an empty array. It then invokes two UNWIND clauses, which iterate over each element in the operated array. For each iterated element, the UNWIND executes the subsequent clause under the context of this element. For example, the array [x] used by the second UNWIND references the variable x, which is 0 when the first UNWIND iterates over the item 0 in the array [0]. After the CALL, the query utilizes FOREACH, whose execution depends on the query context of the subquery in the CALL. To optimize the query execution, Neo4j tries to flatten the FOREACH loop. However, such optimization does not work well when the loop involves update operations (i.e., MERGE) under complicated contexts (i.e., the contexts produced by CALL). The improper optimization corrupts the internal data structures of Neo4j, Eagers, which are the production of another optimization, Eagerness analysis. In the end, an internal error is triggered when Neo4j tries to access the corrupted Eagers. To fix this bug, Neo4j developers modify both the Eagerness analysis and the flattening strategy for FOREACH clauses to ensure they work consistently.
Empirical Comparison. To empirically demonstrate the effectiveness of Dinkel on code coverage and bug detection over the state-of-the-art, we evaluated Dinkel and existing Cypher testing tools (i.e., GDSmith, GDBMeter, GraphGenie, GAMERA, and GRev) on Neo4j and RedisGraph, which are the only two GDBMSs supported by all these tools. Each evaluation persisted for 48 hours. Table 11 shows the comparison results of Dinkel and existing tools.
Table 11. Results of comparison
@@ 6 RETURN n2 AS n3
Covered Line Bug Covered Line Bug
GDsmith 533k 0 15k 0
GDBMeter 283k 1 19k 1
GraphGenie 506k 0 3k 0
GAMERA 283k 1 3k 0
GRev 601k 1 14k 1
Dinkel 737k 11 20k 18
Figure 17. Covered lines of Neo4j and RedisGrpah.
Code Coverage. As shown in Table 11, on average, Dinkel can cover 67% and 85% more code than existing approaches in Neo4j and RedisGraph, respectively. Among the tools except for Dinkel, GRev covers the most code (i.e., 601k lines) in Neo4j. Compared to GRev, Dinkel covers 23% more code. Benefiting from the powerful capability of generating complex Cypher queries, Dinkel can cover much deeper logic of GDBMSs that is related to processing advanced Cypher features and complicated data dependencies. In contrast, simple queries generated by existing tools have little chance to touch such logic. In RedisGraph, except for Dinkel, GDBMeter covers the most code (i.e., 19k lines). Compared to GDBMeter, Dinkel covers 5% more code. The coverage improvement is less significant than in Neo4j, because RedisGraph supports fewer Cypher features and cannot handle some complicated Cypher semantics. For example, RedisGraph does not support processing the subqueries in FOREACH and CALL clauses. As a result, the space for Dinkel to improve the code coverage in RedisGraph is limited. Figure 17 shows the trends of the covered code in each GDBMS during testing. Dinkel and all the state-of-the-art cover new codes quickly and saturate in earlier stages. In almost all testing campaigns, Dinkel can cover more code than the existing tools.
Found Bugs. As shown in Table 11, compared to existing approaches, Dinkel finds more bugs in both Neo4j and RedisGraph. We describe the relations of the found bugs in Figure 18. Note that none of these approaches have found logic bugs within 48 hours, so the bugs shown are related to internal errors or crashes. Specifically, Dinkel found 11 bugs in Neo4j. GDBMeter, GRev, and GAMERA found 1 bug, and this bug can also be found by Dinkel. In RedisGraph, Dinkel found 17 bugs, while GRev and GDBMeter found only 1 bug, and this bug can also be found by Dinkel. For the 27 bugs missed by existing approaches, their bug-triggering queries contain either complicated data dependencies or advanced Cypher features that are not supported by existing approaches. The bug shown in Figure 16 is one of the 27 bugs. These results demonstrate that Dinkel can find more bugs than existing approaches by generating more complex Cypher queries.
Figure 18. Relation of bugs found by existing tools.
@@ 6. Related Work
GDBMS Testing. GDBMS testing is an emerging research field and several approaches (Jiang et al., 2023b; Kamm et al., 2023; Hua et al., 2023; Zheng et al., 2022; Zhuang et al., 2023; Mang et al., 2024) are proposed. All these approaches integrate techniques for constructing test oracles to identify bugs, especially logic bugs (i.e., the bugs cause GDBMSs to return incorrect results). Grand (Zheng et al., 2022) targets GDBMSs using the Gremlin query language and detects logic bugs using differential testing (McKeeman, 1998). Several approaches (Jiang et al., 2023b; Kamm et al., 2023; Hua et al., 2023; Zhuang et al., 2023; Mang et al., 2024) support GDBMSs using Cypher. Similar to Grand, GDsmith (Hua et al., 2023) also uses differential testing to find logic bugs. GDBMeter (Kamm et al., 2023) leverages the idea of query partitioning (Rigger and Su, 2020b), which decomposes the predicate of a query and checks whether the queries with the decomposed predicates produce consistent results with the original query. GraphGenie (Jiang et al., 2023b) modifies the graph patterns used in a query and checks whether the query with modified graph patterns satisfies the expected relationship (i.e., equivalence, subset, or superset) with the original queries. GAMERA (Zhuang et al., 2023) extracts metamorphic relations (Chen et al., 1998) from the manipulated graph data and validates GDBMSs by checking whether the outputs of generated queries satisfy these relations.
Different from all these existing approaches, which focus on new test oracles for finding logic bugs, Dinkel focuses on improving query generation, which is the groundwork of GDBMS testing. An interesting future work is to combine the techniques of Dinkel and existing approaches, enhancing both query generation and test-oracle construction for finding more bugs in GDBMSs.
RDBMS Testing. Compared to GDBMS testing, testing relational database management systems (RDBMSs) is more mature, where both query generation (SQLsmith: A random SQL query generator., 2023; Zhong et al., 2020; Fu et al., 2022; Jiang et al., 2023a; Liang et al., 2023) and test-oracle construction (Rigger and Su, 2020c, a, b; Hao et al., 2023; Song et al., 2023; Tang et al., 2023; Jiang and Su, 2024) are well-researched. We focus on discussing the work on query generation for RDBMSs. SQLsmith (SQLsmith: A random SQL query generator., 2023) embeds the SQL grammar (SQL standard., 1992) and can generate queries containing complex statements. SQUIRREL (Zhong et al., 2020) integrates a new intermediate representation (IR) for modeling the dependencies among statements in a SQL query. Therefore, SQUIRREL can generate queries containing multiple statements. DynSQL (Jiang et al., 2023a) incrementally generates SQL statements for a query. For each statement, DynSQL queries the tested RDBMS to capture the latest DBMS state. In this way, DynSQL can generate complex and valid queries with multiple statements. LEGO (Liang et al., 2023) proposes type-affinity to describe the pattern of composing two types of SQL statements. Based on the type-affinity extracted from existing queries, LEGO can mutate and synthesize new queries that have a higher chance of increasing the code coverage of RDBMSs.
Different from these approaches, Dinkel is designed for GDBMSs and integrates techniques for improving Cypher query generation.
@@
State-Aware Fuzzing. Fuzzing is a promising technique for detecting bugs in various software (Jiang et al., 2020, 2022; American Fuzzy Lop., 2021; Lyu et al., 2019; Yun et al., 2018; Padhye et al., 2019; Chen and Chen, 2018). Some approaches (Ma et al., 2023; Kim et al., 2022; Jiang et al., 2023a; Zhao et al., 2022; Atlidakis et al., 2019) have been proposed to find bugs more efficiently in state-sensitive systems. RESTler (Atlidakis et al., 2019) analyzes the API specification of the tested cloud service and generates request sequences that follow the inferred producer-consumer dependencies. RESTler also collects the response observed during prior executed requests to guide subsequent request sequence generation. LOKI (Ma et al., 2023) proposed to test blockchain consensus protocols. It builds a state model to dynamically track the state transition of each node in the blockchain systems and accordingly generates inputs with proper targets, types, and contents. StateFuzz (Zhao et al., 2022) is designed to test Linux drivers. It utilizes static analysis to recognize critical variables that affect the control flows or memory accesses, and represents program states using these variables. StateFuzz prioritizes test cases triggering new states using three-dimensional feedback. To effectively test USB gadget stacks, FuzzUSB (Kim et al., 2022) extracts the internal state machines from USB gadget drivers via static analysis and symbolic execution, and uses such state information as fuzzing feedback to guide test-case generation.
Similar to these approaches, Dinkel also leverages state information and is the first state-aware testing approach designed for detecting bugs in GDBMSs.
@@ 7. Conclusion
In this paper, we have presented a novel, practical approach for testing GDBMSs using complex Cypher queries. Our approach models the graph state into two categories: query context and graph schema. Based on the model, we have proposed state-aware query generation, which incrementally constructs the clauses for the generated queries and intermittently updates the state information after each clause is constructed. We have implemented our approach as an automatic GDBMS testing framework, Dinkel, and evaluated it on three popular GDBMSs. The evaluation results demonstrate that Dinkel can effectively generate complex Cypher queries with high validity. Dinkel found 60 bugs, many of which had been missed by existing approaches. Considering its effectiveness in Cypher query generation, we believe Dinkel can be the foundation for GDBMS testing, facilitating and inspiring follow-up research, like Csmith (Yang et al., 2011) for compiler testing.

### v3

@@ 5.5. Comparison
Bug Latency Study. To demonstrate that Dinkel can find bugs missed by existing approaches, we investigate the latencies of bugs found by Dinkel. For each bug, we check whether its bug-inducing commit was created before the years when existing approaches were published. If Dinkel finds some long-latent bugs, we can conclude that Dinkel can find bugs missed by existing approaches. This comparison is reasonable and objective because: (1) none of the bugs found by Dinkel are marked as duplicated by developers, meaning that no approach found these bugs until Dinkel found them; and (2) all existing approaches have extensively tested Neo4j and RedisGraph (Kamm et al., 2023; Hua et al., 2023; Jiang et al., 2023b; Zhuang et al., 2023; Mang et al., 2024b), meaning that in these two GDBMSs, no approach found the long-latent bugs found by Dinkel during their evaluation. Some existing approaches do not test Memgraph, but we still include it to show the effectiveness of Dinkel.
GDsmith (Hua et al., 2023), GDBMeter (Kamm et al., 2023), GraphGenie (Jiang et al., 2023b), and GAMERA (Zhuang et al., 2023) were published in 2023, and GRev (Mang et al., 2024b) was published in 2024. Therefore, for each bug, we checked whether its bug-inducing commit was created before 2024 (i.e., in 2023 or earlier) and 2023. As shown in Table 11, among the 113 bugs in Neo4j and RedisGraph, 47 already existed before 2023, among which 10 are logic bugs. It indicates that all existing approaches failed to find these 47 bugs in their extensive evaluation. In addition, 45 bugs (24 logic bugs) were introduced to Neo4j and RedisGraph between 2023 and 2024, while GRev, which was proposed after 2024, cannot find these bugs. In addition, GDsmith and GRev support Memgraph, but neither can find the 4 long-latent bugs induced before 2023. These results indicate that existing approaches indeed miss many long-latent bugs, while Dinkel can effectively find them.
Table 11. Latency of the bugs found by Dinkel. The number in the parentheses is for the logic bugs
@@
Neo4j 51 (13) 40 (9) 27 (7) 2016
RedisGraph 62 (18) 52 (15) 20 (3) 2020
Memgraph 14 (2) 9 (2) 4 (0) 2021
Total 127 (33) 101 (26) 51 (10) 2016
Table 12. The clauses used by bug-triggering queries. The number in the parentheses is for the logic bugs. N/A indicates that the GDBMS does not support the clause
Clause Neo4j RedisGraph Memgraph Total
MATCH 20 (7) 28 (6) 5 (1) 53 (14)
CREATE 17 (9) 26 (11) 2 (1) 55 (21)
MERGE 12 (4) 31 (12) 5 (1) 48 (17)
DELETE 4 (4) 16 (6) 1 (1) 21 (11)
REMOVE 0 1 (0) 0 1 (0)
SET 4 (2) 3 (1) 0 7 (3)
UNWIND 9 (1) 11 (6) 1 (0) 21 (7)
WITH 16 (3) 23 (8) 0 39 (11)
RETURN 40 (8) 39 (13) 11 (1) 90 (22)
CALL 15 (3) 22 (8) 4 (0) 41 (11)
FOREACH 8 (5) 6 (3) 0 14 (8)
UNION 10 (3) 7 (3) 1 (0) 18 (6)
EXISTS 10 (3) N/A N/A 10 (3)
COUNT 10 (4) N/A N/A 10 (4)
Clause Analysis. We analyze the clauses used in the 127 bug-triggering queries and show the results in Table 12. The results indicate that the bugs found by Dinkel are related to various Cypher clauses. Combining Table 4 and Table 12, we can demonstrate that many bugs found by Dinkel cannot be found by existing approaches, because they do not even support the clauses used in some bug-triggering queries (e.g., 14 bugs (8 logic bugs) related to FOREACH clauses). As discussed in Section 4, existing approaches cannot support these clauses because they lack a systematic model for query generation and a general test oracle for query validation.
Empirical Comparison. To empirically demonstrate that Dinkel outperforms the state-of-the-art on code coverage and bug detection, we further evaluated Dinkel and existing approaches on Neo4j and RedisGraph, which are the only two GDBMSs supported by all these approaches. Each evaluation persisted for 48 hours and was repeated 5 times. Table 13 shows the comparison results.
Table 13. Results of covered lines (average $\pm$ standard deviation) and found bugs of existing approaches over 5 runs
@@
Line (Avg $\pm$ SD) Bug Line (Avg $\pm$ SD) Bug
Dinkel 750k $\pm$ 5.2k 19 20k $\pm$ 0.1k 26
GDBMeter 283k $\pm$ 0.2k 1 19k $\pm$ 0.3k 1
GRev 597k $\pm$ 4.0k 1 14k $\pm$ 0.1k 1
GAMERA 283k $\pm$ 0.1k 1 3k $\pm$ 0.1k 0
GDsmith 515k $\pm$ 8.6k 0 15k $\pm$ 0.1k 0
GraphGenie 506k $\pm$ 2.0k 0 3k $\pm$ 0.1k 0
Code Coverage. As shown in Table 13, on average, Dinkel covers 71% and 85% more code than existing approaches in Neo4j and RedisGraph, respectively. Among the tools, excluding Dinkel, GRev covers the most code (i.e., 597k lines) in Neo4j. Compared to GRev, Dinkel covers 25% more code. Benefiting from the powerful query generation, Dinkel can cover much deeper logic of GDBMSs that is related to processing advanced Cypher features and complicated data dependencies. In contrast, simple queries generated by existing tools rarely trigger such logic. In RedisGraph, excluding Dinkel, GDBMeter covers the most code (i.e., 19k lines). Compared to it, Dinkel covers 5% more code. The coverage improvement is less significant than in Neo4j, because RedisGraph supports fewer Cypher features and cannot handle some complicated Cypher semantics. For example, RedisGraph does not support processing the subqueries in FOREACH and CALL clauses. As a result, the space for Dinkel to improve the code coverage in RedisGraph is limited.
Found Bugs. Table 13 shows the number of bugs found by each approach over 5 runs. Figure 12 describes the bug relation. Dinkel finds the most bugs in both Neo4j (19 bugs) and RedisGraph (26 bugs), where logic bugs comprise 4 of the bugs found in Neo4j and 7 in RedisGraph. In Neo4j, GDBMeter, GRev, and GAMERA found 1 bug. In RedisGraph, GRev and GDBMeter found 1 bug. Both of these bugs were also identified by Dinkel. For the 43 bugs missed by existing approaches, their bug-triggering queries contain either complicated data dependencies or advanced Cypher features that are not supported by existing approaches. The bugs shown in Figure 10 and Figure 11 are two of the 43 bugs. These results demonstrate that Dinkel can find more bugs than existing approaches by generating and validating complex queries.
Figure 12. Relation of bugs found by existing approaches.
@@ 6. Related Work
GDBMS Testing. GDBMS testing is an emerging research field, where several approaches (Jiang et al., 2023b; Kamm et al., 2023; Hua et al., 2023; Zheng et al., 2022; Zhuang et al., 2023; Mang et al., 2024b) have been proposed. All these approaches aim to construct test oracles to identify logic bugs. Grand (Zheng et al., 2022) targets GDBMSs using the Gremlin query language and detects logic bugs using differential testing (McKeeman, 1998). Several approaches (Jiang et al., 2023b; Kamm et al., 2023; Hua et al., 2023; Zhuang et al., 2023; Mang et al., 2024b) support GDBMSs using Cypher. Similar to Grand, GDsmith (Hua et al., 2023) also uses differential testing to find logic bugs. GDBMeter (Kamm et al., 2023) leverages query partitioning (Rigger and Su, 2020b), which decomposes the predicate of a query and checks if the queries with the decomposed predicates produce consistent results with the original query. GraphGenie (Jiang et al., 2023b) modifies the graph patterns used in a query and checks if the query with modified graph patterns satisfies the expected relationship (i.e., equivalence, subset, or superset) with the original queries. GAMERA (Zhuang et al., 2023) extracts metamorphic relations (Chen et al., 1998) from the manipulated graph data and checks whether the outputs of generated queries satisfy these relations.
Dinkel is different from all existing approaches in two key aspects. First, Dinkel significantly improves query generation, the core foundation of GDBMS testing. Second, while existing approaches support correctness testing for only queries following specific patterns, Dinkel generally tackles the problems of test-oracle construction, enabling validation for arbitrary queries.
RDBMS Testing. Compared to GDBMS testing, testing relational database management systems (RDBMSs) is more mature, where both query generation (SQLsmith: A random SQL query generator., 2023; Zhong et al., 2020; Fu et al., 2022; Jiang et al., 2023a; Liang et al., 2023) and test-oracle construction (Rigger and Su, 2020c, a, b; Hao et al., 2023; Song et al., 2023; Tang et al., 2023; Jiang and Su, 2024) are well-researched. SQLsmith (SQLsmith: A random SQL query generator., 2023) embeds the SQL grammar (SQL standard., 1992) and can generate complex SQL statements. DynSQL (Jiang et al., 2023a) incrementally generates SQL statements for a query by querying the latest DBMS schema, which enables DynSQL to decouple the generation of each statement and helps it generate queries containing multiple complex statements. To find logic bugs in RDBMSs, PQS (Rigger and Su, 2020c) synthesizes customized queries that fetch specific rows of tables. If the tested RDBMS fails to fetch the rows, a logic bug is identified. Pinolo (Hao et al., 2023) modifies the predicate of a query, where the modified query should produce the subset/superset of the results of the original query. Otherwise, Pinolo identifies a logic bug. EET (Jiang and Su, 2024) transforms the expressions of queries in a semantic-preserving manner and checks whether the queries with the transformed expressions produce the same results as the original queries. EET identifies a bug if their results are not the same. Due to the separation of querying and data manipulation in the SQL language, EET cannot perform clause-level transformations the same way Dinkel can for Cypher queries.
Unlike these approaches for SQL queries, Dinkel is designed for validating Cypher queries. Leveraging the features of the Cypher query language, Dinkel integrates novel techniques for both query generation and test-oracle construction.
@@
State-Aware Fuzzing. Fuzzing is a promising technique for finding bugs in software (Jiang et al., 2020, 2022; American Fuzzy Lop., 2021; Lyu et al., 2019; Yun et al., 2018; Chen and Chen, 2018). Some approaches (Ma et al., 2023; Kim et al., 2022; Jiang et al., 2023a; Zhao et al., 2022; Atlidakis et al., 2019) have been proposed to find bugs more efficiently in state-sensitive systems. RESTler (Atlidakis et al., 2019) analyzes API specifications of the tested cloud services and generates request sequences that follow the inferred producer-consumer dependencies. RESTler also collects the response observed during prior requests to guide subsequent request generation. LOKI (Ma et al., 2023) proposed to test blockchain consensus protocols. It builds a state model to dynamically track the state transition of each node in the blockchain systems and accordingly generates inputs with proper targets, types, and contents. StateFuzz (Zhao et al., 2022) is designed to test Linux drivers. It utilizes static analysis to recognize critical variables that affect control flows or memory accesses, and represents program states using these variables. StateFuzz prioritizes test cases that trigger new states. For effectively testing USB gadget stacks, FuzzUSB (Kim et al., 2022) extracts the internal state machines from USB gadget drivers via static analysis and symbolic execution, before using this state information as fuzzing feedback for guiding test-case generation.
Different from these approaches, Dinkel models GDBMS-related state information for allowing query generation to introduce complex data dependencies, all while retaining query validity.
@@ 7. Conclusion
We have presented a novel and practical framework, Dinkel, for validating GDBMSs. We model the graph state as query context and graph schema, and propose state-aware query generation to generate complex and valid Cypher queries for testing the deep logic of GDBMSs. Moreover, we propose two fine-grained query transformations: clause-level transformations and expression-level transformations, which can operate on arbitrary Cypher queries to validate their correctness. In our evaluation, Dinkel found 127 bugs in three well-known GDBMSs, among which 33 are logic bugs. Considering its significant advancement in bug finding, we believe Dinkel can lay a practical foundation for GDBMS testing, facilitating and inspiring follow-up research on GDBMS reliability.